The perimeter of a triangle is 49 centimeters. The longest side is 5cm less than the sum of the other two sides. Twice the shortest side is 10cm less than the longest side. Find the lengths of each side of the triangle.
Then lengths of the sides, from shortest to longest are..

Answer by stanbon(75887)

The perimeter of a triangle is 49 centimeters. The longest side is 5cm less than the sum of the other two sides. Twice the shortest side is 10cm less than the longest side. Find the lengths of each side of the triangle.
------------------------
Let the shortest side be "x".
---
longest side = 2x + 10
------
2x+10 = x + (midside)-5
So, midside = x+15
-------
Equation:
x + x+15 + 2x+10 = 49
4x = 24
----------------------------
shortest = x = 6 cm
middle = x+15 = 21 cm
longest = 2x+10 = 22 cm
=========================
Cheers,
Stan H.